# Vacancy-mediated transport and segregation tendencies of solutes in FCC nickel under diffusional creep: A density functional theory study

### v1

@@ Abstract
density functional theory, diffusional creep, segregation tendency, nickel alloys, kinetic cluster expansion, Onsager transport coefficients
preprint: APS/123-QED
@@ I Introduction
Atomistic modeling techniques, such as density functional theory (DFT), have primarily been brought to bear for calculating the energetics of point defect diffusion, such as migration, binding, and formation energies of vacancies. Defect energetics obtained from DFT can be incorporated into higher length scale tools, such as kinetic Monte Carlo (KMC) [19] or self-consistent mean field theory (SCMF)[20, 21]. Such tools can compute the total set of Onsager transport coefficients ${L_{ij}}^{\prime}s$. While there are experimental methods to determine ${L_{ij}}$’s, these are generally limited to high temperatures and select compositions. Experimental tracer diffusion coefficients have been used to obtain ${L_{ij}}$’s using relations such as the Darken or Manning model [22]. However, such models cannot predict negative values for off-diagonal components of ${L_{ij}}$ [23, 24]. The accurate determination of Onsager transport coefficients is useful not only in determining vacancy and solute diffusivities via the diagonal components but also in the analysis of flux coupling of solute species with vacancies via the off-diagonal components. Furthermore, the total set of Onsager transport coefficients is essential to accurately parameterize mean-field and rate-theory models for diffusion creep, thermal nonequilibrium segregation, and radiation-induced segregation  [25, 26, 27].
@@
In this work, we employ DFT to obtain the vacancy-mediated diffusion energetics of Cr, Re, Ta, W, Mo, and Co in FCC Ni. By providing the DFT energetics as input into the SCMF framework implemented in the KineCluE code, we evaluate the total set of Onsager transport coefficients, $L_{ij}$’s, in the limit of dilute solute concentrations. The Onsager coefficients are then employed to investigate the tendency of solutes for nonequilibrium segregation at vacancy sinks, such as GBs. Furthermore, using the linear elastic approach implemented in KineCluE, we incorporate the effects of tensile hydrostatic strain on the Onsager transport coefficients to obtain further insights into the transport mechanisms and segregation tendencies. These results are then discussed in the context of GBs in polycrystalline materials under diffusion creep conditions. In combination with experimental testing in the future, the theoretical predictions from the present work are expected to help verify the hypothesis that nonequilibrium elemental segregation occurs at GBs when the diffusional creep mechanism is active.
@@ II.3 Density Functional Theory Computational Details
DFT calculations are performed using the projector-augmented wave method [46] as implemented in the Vienna ab initio Simulation Package [47, 48]. Spin-polarized calculations with collinear magnetic moments with an initial value of 2 Bohr magnetons ($\mu$B) per atom were implemented to generate a ferromagnetic Ni system, The relaxed FCC Ni structure has a magnetic moment of 0.6 $\mu$B per atom. The same value of 2 $\mu$B was used to initialize the magnetic moment of the solute atoms, Co, Re, Mo, Ta, and W which yielded a final magnetic moment value of 1.8, -0.1, -0.2, -0.2, -0.2 $\mu$B, respectively. The negative sign denotes a spin-down state (opposite to that of the Ni atoms). In the case of Cr solute, initial guesses of spin-up (2 $\mu$B ) and spin-down (–2 $\mu$B ) states were tested but both yielded the same final magnetic moment of 2.1 $\mu$B. The Perdew-Burke-Ernzerhof [49] generalized-gradient approximation (GGA) is adopted for the exchange-correlation functional. A plane-wave cutoff energy of 450 eV was set in all calculations. The convergence criteria for total energies and forces were set to $10^{-6}$ eV and 0.005 eV/Å, respectively, for relaxing defect-containing supercells at constant volume. For climbing-image nudged elastic band calculations [50, 51, 52], one intermediate image was used to obtain the saddle point energies along migration paths. The convergence criteria in climbing-image nudged elastic band calculations were set to $10^{-5}$ eV and 0.01 eV/Å, for total energies and forces, respectively. $3\times 3\times 3$ (108 atoms) and $4\times 4\times 3$ (192 atoms) supercells were used for calculating point defect formation, binding, and migration energies. $5\times 5\times 5$ and $4\times 4\times 5$ k-point meshes were used for the Brillouin zone sampling of the 108-atom and 192-atom supercells, respectively. The results from both supercell sizes agree well for all solute species, except for Cr-vacancy pairs where the 108-atom supercell gives overestimated migration barriers for certain vacancy jumps (especially the $1nn\rightarrow 1nn$ jump). The migration barriers obtained for the Ni-Cr system from the 192-atom supercell have better agreement with the results reported by Toijer et al. [37] using 256-atom supercells. This gives us confidence in our 192-atom results. The results presented in this work are based on the 192-atom supercell DFT data unless noted otherwise.
@@ III.1 Density Functional Theory Results
Figure 2: Solute-vacancy binding energies as a function of nearest neighbor distance in FCC nickel. Positive binding energies are attractive and negative binding energies are repulsive.
@@ III.2 Solute Diffusion Coefficients
The solute diffusivities calculated using this approach are valid for dilute concentrations of $[B]<0.094\%$ (see Appendix A). The diffusivities of the six solutes are plotted versus inverse temperature in Figure 4. Experimental tracer diffusivity data from Refs. [64, 65, 66, 67, 68, 69, 70] are plotted alongside for comparison. The results show a reasonable agreement between the KineCluE calculations and the experimental data. The calculated diffusivities were fit to the Arrhenius equation $D=D_{0}e^{-Q/{k_{B}T}}$ to determine the activation energies $Q$ and the prefactors $D_{0}$, which are listed in Table 3. We also performed these calculations for the solvent Ni. The activation energy for Ni self-diffusion was obtained as 2.51 eV (i.e., Q = $E_{f}^{V}+E_{m}^{V}$ = 1.42 eV + 1.09 eV). Amongst the solutes, Re, W, and Co (the slow diffusers) yielded higher effective activation energies of 2.97, 2.71, and 2.60 eV, respectively, while Mo, Cr, and Ta (the fast diffusers) yielded low activation energies of 2.47, 2.44, and 2.36 eV, respectively.
@@ III.5 Relevance to Diffusional Creep
Considering strain effects, the results of Section III.4, which pertain to tensile strain, are relevant to the perpendicular boundaries. However, the segregation tendencies reported therein must be reversed since the results were interpreted for segregation at vacancy sinks, whereas perpendicular boundaries act as vacancy sources. For parallel GBs (vacancy sinks), the effects of compressive strain are more relevant. Without a loss of generality, we may extrapolate the trends in $G_{V}$ and $D_{pd}$ obtained for tensile strains towards compressive strains. Such an extrapolation of our results indicates that compressive strain in the vicinity of parallel GBs will lower $T_{drag}$, thus making IKE the dominant mechanism. This is expected to strengthen the solute segregation tendencies predicted at these GBs in the absence of strain. Hence, a stronger depletion of Cr, Mo, and Ta and a stronger enrichment of Re and W are expected. Additionally, strain effects are expected to have little influence on the segregation tendency of Co. The results provide us a basis for choosing suitable alloy systems and the expected solute segregation tendencies at GBs for our future experimental investigations to verify the occurrence of diffusional creep mechanisms.
@@ Appendix B DFT Data
Table 7: Vacancy migration barriers in eV for atomic the vacancy jumps in Ni-X systems. X: Cr, Re, Ta, W, Mo, Co. Values from previous DFT studies in the literature [37, 32] are shown in parentheses for comparison

### v2

@@ Abstract
preprint: Accepted copy of paper published in Physical Review Materials. https://doi.org/10.1103/PhysRevMaterials.8.083605
@@ I Introduction
Atomistic modeling techniques, such as density functional theory (DFT), have primarily been brought to bear for calculating the energetics of point defect diffusion, such as migration, binding, and formation energies of vacancies. Defect energetics obtained from DFT can be incorporated into higher length scale tools, such as kinetic Monte Carlo (KMC) [19] or self-consistent mean field theory (SCMF)[20, 21]. Such tools can compute the total set of Onsager transport coefficients ${L_{ij}}$. While there are experimental methods to determine ${L_{ij}}$, these are generally limited to high temperatures and select compositions. Experimental tracer diffusion coefficients have been used to obtain ${L_{ij}}$ using relations such as the Darken or Manning model [22]. However, such models cannot predict negative values for off-diagonal components of ${L_{ij}}$ [23, 24]. The accurate determination of Onsager transport coefficients is useful not only in determining vacancy and solute diffusivities via the diagonal components but also in the analysis of flux coupling of solute species with vacancies via the off-diagonal components. Furthermore, the total set of Onsager transport coefficients is essential to accurately parameterize mean-field and rate-theory models for diffusion creep, thermal nonequilibrium segregation, and radiation-induced segregation  [25, 26, 27].
@@
In this work, we employ DFT to obtain the vacancy-mediated diffusion energetics of Cr, Re, Ta, W, Mo, and Co in FCC Ni. By providing the DFT energetics as input into the SCMF framework implemented in the KineCluE code, we evaluate the total set of Onsager transport coefficients, $L_{ij}$, in the limit of dilute solute concentrations. The Onsager coefficients are then employed to investigate the tendency of solutes for nonequilibrium segregation at vacancy sinks, such as GBs. Furthermore, using the linear elastic approach implemented in KineCluE, we incorporate the effects of tensile hydrostatic strain on the Onsager transport coefficients to obtain further insights into the transport mechanisms and segregation tendencies. These results are then discussed in the context of GBs in polycrystalline materials under diffusion creep conditions. In combination with experimental testing in the future, the theoretical predictions from the present work are expected to help verify the hypothesis that nonequilibrium elemental segregation occurs at GBs when the diffusional creep mechanism is active.
@@ II.3 Density Functional Theory Computational Details
DFT calculations are performed using the projector-augmented wave method [46] as implemented in the Vienna ab initio Simulation Package [47, 48]. Spin-polarized calculations with collinear magnetic moments with an initial value of 2 Bohr magnetons ($\mu$B) per atom were implemented to generate a ferromagnetic Ni system, The relaxed FCC Ni structure has a magnetic moment of 0.6 $\mu$B per atom. The same value of 2 $\mu$B was used to initialize the magnetic moment of the solute atoms, Co, Re, Mo, Ta, and W which yielded a final magnetic moment value of 1.8, -0.1, -0.2, -0.2, -0.2 $\mu$B, respectively. The negative sign denotes a spin-down state (opposite to that of the Ni atoms). In the case of Cr solute, initial guesses of spin-up (2 $\mu$B ) and spin-down (–2 $\mu$B ) states were tested but both yielded the same final magnetic moment of 2.1 $\mu$B. The Perdew-Burke-Ernzerhof [49] generalized-gradient approximation (GGA) is adopted for the exchange-correlation functional. A plane-wave cutoff energy of 450 eV was set in all calculations. The convergence criteria for total energies and forces were set to $10^{-6}$ eV and 0.005 eV/Å, respectively, for relaxing defect-containing supercells at constant volume. For climbing-image nudged elastic band (CI-NEB) calculations [50, 51, 52], one intermediate image was used to obtain the saddle point energies along migration paths. Typically, a single intermediate image is sufficient for a vacancy jump; this provides fast convergence of the saddle point energies and is commonly used in the literature [53, 54]. Additionally, three-image CI-NEB calculations were also tested for the Cr and Re solute-vacancy exchange jumps. The obtained saddle points agree with the single-image calculations to within 1 meV. The convergence criteria in CI-NEB calculations were set to $10^{-5}$ eV and 0.01 eV/Å, for total energies and forces, respectively. $3\times 3\times 3$ (108 atoms) and $4\times 4\times 3$ (192 atoms) supercells were used for calculating point defect formation, binding, and migration energies. $5\times 5\times 5$ and $4\times 4\times 5$ k-point meshes were used for the Brillouin zone sampling of the 108-atom and 192-atom supercells, respectively. The results from both supercell sizes agree well for all solute species, except for Cr-vacancy pairs where the 108-atom supercell gives overestimated migration barriers for certain vacancy jumps (especially the $1nn\rightarrow 1nn$ jump). The migration barriers obtained for the Ni-Cr system from the 192-atom supercell have better agreement with the results reported by Toijer et al. [37] using 256-atom supercells. This gives us confidence in our 192-atom results. The results presented in this work are based on the 192-atom supercell DFT data unless noted otherwise.
@@ III.1 Density Functional Theory Results
Figure 2: Solute-vacancy binding energies as a function of nearest neighbor distance in FCC nickel. The nearest neighbor number refers to the position of vacancy with respect to the solute atom. Positive binding energies are attractive and negative binding energies are repulsive.
@@ III.2 Solute Diffusion Coefficients
The solute diffusivities calculated using this approach are valid for dilute concentrations of $[B]<0.094\%$ (see Appendix A). The diffusivities of the six solutes are plotted versus inverse temperature in Figure 4. Experimental tracer diffusivity data from Refs. [66, 67, 68, 69, 70, 71, 72] are plotted alongside for comparison. The results show a reasonable agreement between the KineCluE calculations and the experimental data. The calculated diffusivities were fit to the Arrhenius equation $D=D_{0}e^{-Q/{k_{B}T}}$ to determine the activation energies $Q$ and the prefactors $D_{0}$, which are listed in Table 3. We also performed these calculations for the solvent Ni. The activation energy for Ni self-diffusion was obtained as 2.51 eV (i.e., $Q=E_{f}^{V}+E_{m}^{V}$ = 1.42 eV + 1.09 eV). Amongst the solutes, Re, W, and Co (the slow diffusers) yielded higher effective activation energies of 2.97, 2.71, and 2.60 eV, respectively, while Mo, Cr, and Ta (the fast diffusers) yielded low activation energies of 2.47, 2.44, and 2.36 eV, respectively. Comparing the values of $Q$ with the migration barriers of solute-vacancy exchange ($E_{m}^{V-B}$) in Table 7 and Figure 3, we find a strong correlation. This is because Q can be roughly estimated by adding $E_{f}^{V}$—which remains constant regardless of the solute—and $E_{m}^{V-B}-E_{b}^{1nn}$ under the assumption of ideal random walk behavior [73]; here, $E_{b}^{1nn}$ is the binding energy of the solute B with a vacancy in the first nearest neighbor position. The above indicates that the solute-vacancy exchange controls the vacancy diffusion mechanism of these solutes and that deviations from random walk behavior due to correlation effects are minimal, especially for Co, Mo, Re, and W. For the two fastest diffusers, Cr and Ta, kinetic correlation effects are more prevalent and the sum of the formation, binding, and migration energies does not accurately equal the exact value of Q; this difference is the activation energy of the correlation factor.
@@ III.5 Relevance to Diffusional Creep
Considering strain effects, the results of Section III.4, which pertain to tensile strain, are relevant to the perpendicular boundaries. However, the segregation tendencies reported therein must be reversed since the results were interpreted for segregation at vacancy sinks, whereas perpendicular boundaries act as vacancy sources. For parallel GBs (vacancy sinks), the effects of compressive strain are more relevant. Without a loss of generality, we may extrapolate the trends in $G_{V}$ and $D_{pd}$ obtained for tensile strains towards compressive strains. Such an extrapolation of our results indicates that compressive strain in the vicinity of parallel GBs will lower $T_{drag}$, thus making IKE the dominant mechanism. This is expected to strengthen the solute segregation tendencies predicted at these GBs in the absence of strain. Hence, a stronger depletion of Cr, Mo, and Ta and a stronger enrichment of Re and W are expected. Additionally, strain effects are expected to have little influence on the segregation tendency of Co. The results provide us a basis for choosing suitable alloy systems and the expected solute segregation tendencies at GBs for our future experimental investigations to verify the occurrence of diffusional creep mechanisms. In a recent mesoscale modeling study, Hussein et al. [78] showed segregation of elements in Fe-Au alloy under creep conditions. The results of their diffusion model coupled with polycrystal elastic stress fields demonstrated that the segregation tendency is affected by the type and extent of stress field around GBs. For instance, Au, the faster solute, was predicted to enrich at the GBs under tensile stress. Our results and methodology enable more accurate parameterization of such mesoscale models.
@@ Appendix B DFT Data
Table 7: Vacancy migration barriers in eV for atomic the vacancy jumps in Ni-X systems. X: Cr, Re, Ta, W, Mo, Co. Values from previous DFT studies in the literature [37, 32] are shown in parentheses for comparison.